OpenStudy (anonymous):

For finding the equation of a straight line do you use y1-y2/x1-x2 or y2-y1/x2-x1?

OpenStudy (anonymous):

the second one

OpenStudy (anonymous):

http://www.purplemath.com/modules/slope.htm according to this site it's the other one?

OpenStudy (accessdenied):

Both are correct.

OpenStudy (anonymous):

can you explain breifly?

OpenStudy (anonymous):

the negatives will cancel out no matter which way you do it

OpenStudy (anonymous):

thank you

OpenStudy (anonymous):

Let me give u an example. consider a straight line y= x 2 points on the line will be (1,1) and (2,2) y1-y2/x1-x2 = 1-2/1-2 = -1/-1 = 1 y2-y1/ x2-x1 = 2-1/2-1= 1/1= 1

OpenStudy (phi):

this explains it pretyy well http://www.khanacademy.org/video/slope-of-a-line?topic=core-algebra

OpenStudy (anonymous):

yeah they work out the same... thanks. Are you familiar with boolean algebra?

OpenStudy (phi):

post your specific question and I'm sure someone will answer.

OpenStudy (anonymous):

i did 20 min ago

OpenStudy (phi):

b(1+a'd) + a(b'c +c'd+d') = b+a(b'c+c'd+d') =? to simplify b'c+c'd+d' use DeMorgan's rules not(a+b)= a'b' not(ab)=a'+b' c'd+d' = not( not(c'd)d)= not((c+d')d)= not(cd + d'd)= not(cd)= c'+d' so b'c + c'd+d' -> b'c + c'+ d' similarly b'c + c = not(not(b'c)c')=not((b+c')c)=not(bc+c'c)= not(bc)=b'+c' which makes b'c + c'+ d' -> b'+c'+d' and the original expression is b+a(b'c+c'd+d') = b+a(b'+c'+d')= b+ab' + a(c'+d') simplify b+ab': b+ab'= not(b'(not(ab'))=not(b'(a'+b))= not(a'b'+b'b)=not(a'b')=a+b we now have b+ab' + a(c'+d') -> a+b+a(c'+d')= b+a(1+(c'+d'))=b+a

OpenStudy (anonymous):

wow thanks a lot

OpenStudy (phi):

let me fix this line (in case you go through this carefully) similarly b'c + c = not(not(b'c)c')=not((b+c')c)=not(bc+c'c)= not(bc)=b'+c' should be b'c + c' = not(not(b'c)c)=not((b+c')c)=not(bc+c'c)= not(bc)=b'+c'